OpenStudy (anonymous):

Find the interval in which k shuld lie so that the roots of the equation 2x^2-kx + 8 =0 are b/w -1 and 4

OpenStudy (anonymous):

(x-a) (x-b)=0 x^2 -(a+b) x + ab =0 2x^2 - 2(a+b) x + 2ab =0 => 2(a+b) =k and => 2ab=8

OpenStudy (anonymous):

I think..i Did nt get u...

OpenStudy (anonymous):

Let a and b be the roots of 2x^2-kx + 8 =0

OpenStudy (anonymous):

Now u got it?

OpenStudy (anonymous):

Ok...So..Hw to Find the Interval

OpenStudy (anonymous):

=>2ab =8 (-1<=a<=4 and -1<=b<=4)

OpenStudy (anonymous):

=> 2(a+b) =k 4(a+b)^2 =k^2 4a^2 + 8ab + 4b^2 =k^2 4a^2 + 24 + 4b^2 =k^2 k^2 =4a^2 +4b^2 +24

OpenStudy (anonymous):

Now find the max value of k and minimum value of k

OpenStudy (experimentx):

 play around these.